Treatment of a 0.25000-g sample of impure potassium chloride with an excess of AgNO3 resultion of 0.2912g of AgCl. Calculate the percentage of CCl in the sample.

1
Expert's answer
2022-05-26T09:22:04-0400

Solution:

Calculate the moles of AgCl:

The molar mass of AgCl is 143.32 g/mol

Therefore,

Moles of AgCl = (0.2912 g AgCl) × (1 mol AgCl / 143.32 g AgCl) = 0.0020318 mol AgCl


Balanced chemical equation:

KCl(aq) + AgNO3(aq) → AgCl(s) + KNO3(aq)

According to stoichiometry:

1 mol of KCl produces 1 mol of AgCl

mol of KCl produces 0.0020318 mol of AgCl

Thus,

Moles of KCl = X = (0.0020318 mol AgCl) × (1 mol KCl / 1 mol AgCl) = 0.0020318 mol KCl


Calculate the mass of KCl:

The molar mass of KCl is 74.55 g/mol

Therefore,

Mass of KCl = (0.002032 mol KCl) × (74.55 g KCl / 1 mol KCl) = 0.15147 g KCl


Calculate the percentage of KCl in the sample:

%KCl = (Mass of KCl / Mass of sample) × 100%

%KCl = (0.15147 g / 0.25000 g) × 100% = 60.59%

%KCl = 60.59%


Answer: The percentage of KCl in the sample is 60.59%